OpenStudy (anonymous):

can some one help me solve the equation Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)).

OpenStudy (kagıtucak):

First find g(-1) g(-1)=1+7-9=-1 then it became to find f(-1) -1-2=-3

OpenStudy (blank ):

I did the math. @kagıtucak is correct. It is -3.

OpenStudy (anonymous):

can you show me how to do it

OpenStudy (blank ):

Sure, just a second.

OpenStudy (blank ):

f(x)=x-2 g(x)=x^(2)-7x-9 f(g(-1)) Setup the composite result function. fo+g notation is interpreted as f(g(x)). Evaluate f(g(x)) by substituting in the value of g into f. f(x^(2)-7x-9)=x^(2)-7x-9-2 Subtract 2 from -9 to get -11. f(x^(2)-7x-9)=x^(2)-7x-11 Evaluate the result function by replacing x with -1 in the function. (-1)^(2)-7(-1)-11 Squaring an expression is the same as multiplying the expression by itself 2 times. ((-1)(-1))-7(-1)-11 Multiply -1 by -1 to get 1. (1)-7(-1)-11 Multiply -7 by the -1 inside the parentheses. (1)+7*1-11 Multiply 7 by 1 to get 7. 1+7-11 Add 7 to 1 to get 8. 8-11 Subtract 11 from 8 to get -3. -3 Hope that will help you step by step.

OpenStudy (anonymous):

thank @Blank 

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)